BARRODU143@gmail.com

Grade:12

1 Answers

Aman Bansal
11 years ago

Dear Veeru,

For the electron to travel in a straight line, the net force must be zero. This means that the electric
force equals the magnetic force
q E = qvB
B = E
v
We can now find the field and velocity in terms of given quantities.
V2 = Ed
E = V2
d
= 100V
20 ×10−3m
= 5000V / m
1
2
mv2 = qV1
v = 2qV1
m
=1.87 ×107m / s
B = 5000V / m
1.87 ×107 = 2.675 ×10−4 T

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. 

http://www.askiitians.com/packages/packages.aspx

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert